OpenStudy (anonymous):

What is the solution to the system of equations? x=4y+6 x-3y=4 ( , )

OpenStudy (sleepyhead314):

try plugging (4y+6) for x into x - 3y = 4 can you do that? :) then solve for y

OpenStudy (anonymous):

where?

OpenStudy (sleepyhead314):

(4y+6) for x into -> x - 3y = 4

OpenStudy (anonymous):

(4y+6)x-3y=4

OpenStudy (sleepyhead314):

close :) I meant something like 4y+6 - 3y = 4

OpenStudy (sleepyhead314):

now can you solve for y from that? :)

OpenStudy (anonymous):

y = 2

OpenStudy (sleepyhead314):

close! :D when you subtract 6 to the other side 4 - 6 = ?

OpenStudy (anonymous):

-2

OpenStudy (sleepyhead314):

right! :) that would be y ^_^

OpenStudy (anonymous):

:)

OpenStudy (sleepyhead314):

now lets plug that into x = 4(y) + 6

OpenStudy (anonymous):

x = -2

OpenStudy (sleepyhead314):

that's right ^_^ now the format for a coordinate is (x ,y)

OpenStudy (anonymous):

(-2,-2)

OpenStudy (anonymous):

Thank you very much :D

OpenStudy (sleepyhead314):

Glad I could help! :)